OpenStudy (anonymous):

Write the solution in lowest terms. 1/3 - (-3/4) = 4/12 1/3

OpenStudy (anonymous):

or 1 and 1/12

OpenStudy (anonymous):

13/12

OpenStudy (anonymous):

1/3 - (-3/4) = 1/3 + 3/4 = 4/12 + 9/12 = 13/12